# The connection of polaritonic chemistry with the physics of a spin glass

### v1

@@ 3.1 Collective strong coupling with the cavity Hartree equations
Figure 3: Local polarization features of Shin-Metiu molecules under collective VSC. Left: Finite cavity-induced molecular polarizations (blue) that emerge due to the self-consistency cycles of the cH equations (reproduced from Ref. 20). The collective Rabi splitting was kept fixed, when increasing the ensemble size $N$. In contrast, the macroscopic polarization (black) quickly drops to zero, within numerical errors. Notice that the numerical results of the Shin-Metiu molecule were determined at the onset of a numerical instability (phase transition), which effectively prevents to reach considerably stronger collective coupling strength with the chosen setup. Right: Probability density of the cavity-induced polarizations calculated for $N=1024$ molecules of the above setup. The system was equilibrated for $600$ time steps and then polarization data was considered from another $600$ time steps. Two striking features appear: First, the local polarizations are symmetrically distributed. Second, the probability distribution seems almost discontinuous, even though the electronic-structure data is time-averaged over many vibrational cycles. This should smoothen the data, unless some frustration mechanism counteracts.
3.2 Nuclear magnetic resonance experiments
In the previous subsection, we have seen that solving the cH equations reveals several fascinating physical mechanisms in an ensemble of molecules embedded in an optical cavity at ambient temperature $T$ (polarization ordering, frustration, potential phase transitions). In the following, we compare the above theoretical predictions with recent experimental nuclear magnetic resonance (NMR) results under VSC. 27 In Figs. 4a-c the influence of VSC on the equilibrium concentration between two conformations of a molecular balance sensitive to London dispersion forces is studied with NMR spectroscopy. The experiments reveal that VSC can indeed modify the equilibrium rate constant and thus changes the chemical properties if tuned on resonance with a specific C-H stretching mode of the solute molecules. If we disregard the (important) resonance feature for the moment, the following three experimental observations seem to directly relate to our previous discussions: First, the absence of cavity-induced chemical shifts indicates that the electronic structure is on average not polarized by the cavity, which perfectly agrees with what we would anticipate from Eq. (9). Second, the broadening of the chemical shifts seems modified under VSC, which one would expect for a cavity-induced polarization glass, i.e., from Eq. (10). Notice, however, that in Ref. 27 the modified broadening is assigned to experimental artifacts. Nevertheless, it would be interesting to verify if the predicted broadening caused by $P_{\rm cH}(\Delta\mu_{0})$ contributes to the measured data or if the polarization-glass mechanism is too small to be detected in this NMR setup. The third important insight is an abrupt change of the equilibrium constant at a specific collective strong coupling strength (influenced by the concentration) that does not scale further with the collective Rabi splitting. This suggests a phase transition at a specific collective coupling, as we anticipated from the above discussions about the convergence of the cH equations. Overall, the NMR results of the Ebbesen group seem perfectly in line with the above theoretical description based on the formation of a polarization glass except for one caveat: The observed resonance condition has so far not been studied in detail in the numerical simulations presented above. We will come back to this important aspect at the end of this review (see Sec. 4.2.3), where we will argue that the resonance mechanism is likely a manifestation of the dynamical aspect of the cavity-induced frustration mechanism.
Figure 4: Influence of VSC on the London dispersion-force-driven equilibrium, determined from NMR measurements according to Ref. 27: (a) Free energy difference between folded (1f) and unfolded (1u) conformer in benzened-$d_{6}$. (b) On- (red) and off-resonance (black) 1H NMR spectra of a 1 M solution. The shaded peaks allow to distinguish the two different conformers. The absence of a frequency shift between the two spectra indicates that on average the electronic structure is not affected (no chemical shift). The magnitude of the peaks can be related to the equilibrium constant between the two conformers, which indicates VSC-induced chemical changes. (c) Varying the concentration (i.e., the collective coupling strength) reveals a critical concentration (dotted line), where suddenly a different conformational equilibrium constant is approached, which suggests the emergence of a cavity-induced phase transition.
3.3 Connecting polarization and spin glasses
In the previous sections we have concluded, based on numerical and experimental evidence, that under VSC the cavity-mediated electronic structure seems to form electronic polarization patterns (ordering). Here we want to elucidate these features by comparing to the much better understood physics of spin glasses.
Figure 5: Pictorial comparison of a spin glass and a cavity-mediated polarization glass. Left: Spin-frustrated triangular lattice (sketch modified from Ref. 76), whose physical properties can, for example, be modeled by the Edwards-Anderson model 73, based on (random) short-range interactions between the nearest-neighbour spins on a lattice. Right: A dilute molecular polarization glass formed by vibrational strong coupling in optical cavities.20 Its polarizations can assume continuous values, which are mediated by an all-to-all dipole interactions term instead.
That we are using spin glasses as a guideline to understand these polarization effects can be rationalized even pictorially. In Fig. 5 we compare a sketch of a frustrated spin glass with the sketch of a polarization glass. The interaction between the spins/polarizations leads to an frustration effect, where flipping one spin/polarization will force also the others to adapt. We note that usually models of spin glasses are based on (random) short-ranged interactions between spins on a crystal lattice, e.g., the Edwards-Anderson model 73. However, in the polarization-glass case we have a long-range (all-to-all) interaction. While this seems like a very different situations, the best studied spin-glass model, the Sherrington-Kirkpatrick (SK) model 74, 75, is describing almost exactly this case. There are, however, still several distinctions between the case of polaritonic chemistry and standard spin glass systems (models). Most noteworthy is the fact that a priori molecular polarizations are continuous. But there are several other distinctions, such as the coupling and back-action to other physical degrees of freedom (displacement and nuclear coordinates), which we will highlight throughout this work and which open up novel research questions at the interface of these two fundamental and distinct research communities.
@@ 3.3 Connecting polarization and spin glasses
$H_{J}(\boldsymbol{\sigma})=-\sum_{j<i}^{N}J_{ij}\sigma_{i}\sigma_{j}$ (11)
and the couplings $J_{ij}$ between the spins are independent random variables taken from the normal distribution $\mathscr{N}(J_{0}/N,J^{2}/N)$. The presence of an additional external magnetization field $h$, acting on the spins as $\sum_{i}^{N}h\sigma_{i}$ to generate a finite magnetization $m$, can be recast into a finite mean-value $J_{0}=h/m$ of the random interactions $J_{ij}$77 To understand the basic physical properties of the SK model at finite temperature $T$, we follow Refs. 75, 78 and continue with some definitions. The local (”single molecule”) magnetization at temperature $T$ is given by,
@@
where $\alpha$ denotes a possible quasi-thermal equilibrium state of the SK model, i.e., a local minimum in the phase space for a given choice of $J$, at temperature $T$. In more detail, the magnetizations $m(i)_{\alpha}$ correspond to the $\alpha$-th solution of the mean-field equation, 79, 78
@@
which becomes an exact description for the SK model in the thermodynamic limit. 74, 80, 78 The determination of the exponentially large number of solutions is a computationally very demanding task. Eventually, the thermal equilibrium (”single molecular”) magnetization, averaged over all $\alpha$ and all possible choices of J, is defined as
@@
We note that averaging over all $\alpha$ and $J$ makes the result site-independent, i.e., the left-hand side of Eq. (14) is independent of $i$. The numerous fundamental physical properties of the SK model and its implications on polaritonic chemistry will be discussed extensively throughout Secs. 3.5 and 4.
3.4 Mapping the cavity Hartree (cH) equations on the Sherrington-Kirkpatrick (SK) model of a spin glass
We have noticed that the continuous all-to-all dipole interaction of the cavity Hartree Eq. (7) resembles the discrete random spin-spin interaction of the SK model in Eq. (11). In the following we show an explicit mapping of the electronic cH equations on the SK model of a spin glass for a specific case. For this purpose and without loss of generality, we look at an ensemble of $N$ hydrogen-like molecular ions in three dimensions. The single electron free-space Hamiltonian is given in atomic units as 81
$\hat{H}_{i}^{m}=-\frac{\nabla_{\boldsymbol{r}_{i}}^{2}}{2}-\frac{1}{|\boldsymbol{r}_{1,i}|}-\frac{1}{|\boldsymbol{r}_{2,i}|}+\frac{1}{R_{i}},$ (15)
with $\boldsymbol{r}_{1,i}=\boldsymbol{r}_{i}-\frac{\boldsymbol{R}_{i}}{2}$ and $\boldsymbol{r}_{2,i}=\boldsymbol{r}_{i}+\frac{\boldsymbol{R}_{i}}{2}$. Following the common linear combination of atomic-orbital approach of molecular physics, the single-molecule trial wave function is
$\Psi_{i}(\boldsymbol{r}_{i})$ $=$ $\frac{a_{i}\psi(\boldsymbol{r}_{2,i})+\sigma_{i}b_{i}\psi(\boldsymbol{r}_{1,i})}{\sqrt{N_{i}}},$ (16)

### v2

@@ 2.3.1 Polarization glass
Figure 3: Local polarization features of Shin-Metiu molecules under collective VSC. Finite cavity-induced molecular polarizations (blue) that emerge due to the self-consistency cycles of the cavity Hartree equations (reproduced from Ref. 20). The collective Rabi splitting was kept fixed, when increasing the ensemble size $N$. In contrast, the macroscopic polarization (black) quickly drops to zero, within numerical errors. Notice that the numerical results of the Shin-Metiu molecules were determined at the onset of a numerical instability (phase transition), which effectively prevents to reach considerably stronger collective coupling strength with the chosen setup.
3 Mapping cavity-mediated molecular electron correlations to a spin glass
We have seen that local modifications can survive in the large-$N$ limit even at equilibrium if we go beyond the harmonic approximation. Thus under collective VSC, electronic fluctuations are enhanced when compared to the free-space case. The collective SCF convergence issue suggests that many electronic configurations have roughly the same ensemble energy. Thus one runs into an energetically very dense intermolecular spectral region and hence faces huge (at least numerical) degeneracies. That such a situation appears is to be expected, because on the macroscopic ensemble level, entropic/statistical contributions become important. That is, we need to weigh the energy with its number (density) of states and perform free-energy considerations. In order to do so, we need to get information on the electronic density of states, i.e., we need to keep track of how many excited ensemble states are near to the (potentially highly degenerate) ground state.
In quantum chemistry, there is a simple way to get this information. One can perform a configuration-interaction (CI) singles calculations on top of Hartree-Fock Slater determinants 50, i.e., we include electron exchange and correlations. To do so, we leave the dilute-gas assumption and consider the full cHF problem of Eq. (9). That is, the electronic-structures of the individual molecules can overlap and we consider the full cHF ensemble wave function. We remind the reader that on the intermolecular scale, the cHF wave function is supposed to become asymptotically exact, and hence we can neglect CI doubles and triples for the ensemble problem. Before we continue, with CI singles, we briefly would like to mention that over the past years different post-Hartree-Fock methods were proposed to account for cavity-modified correlation effects. For example, coupled-cluster61, 60, 90, Moeller-Plesset perturbation theory91, 92, 93, 94, configuration interaction,95 multi-configurational96, 97, DFT-based 98, 99, 100, 101, 102, 69, 103 and other correlated methods104, 105, 93 are known. However, typically, ab-initio methods that are designed to accurately capture electron correlations are naturally restricted to only a few molecules. A notable generalization to the collective regime can be found in Ref.106 for a coupled cluster-model, which seems to accurately describe local (e.g. intra-molecular) correlation energies under collective electronic strong coupling in a dilute gas. Alternatively, a radiation-reaction embedding approach was proposed in Ref. 59 or machine learning of forces107. However, the later two approaches sacrifice the collectively-induced20 feed-back effects and thus do not account for collectively induced degeneracies and its consequences on the inter-molecular electron correlations. Instead, our subsequent theoretical framework of a polarization/spin glass is fundamental for describing polaritonic chemistry. The self-consistency is an essential ingredient for the understanding on the local vs. collective interplay of realistic molecular ensembles under collective VSC.
For a given cHF solution $|\Psi_{0}\rangle$ (reference ensemble determinant), the the lowest lying excited states are found by 50
$|\Phi_{\boldsymbol{s}}\rangle=\sum_{c}^{N}\sum_{t}^{\infty}s_{c}^{t}|\Psi_{c}^{t}\rangle,$ (19)
where $\boldsymbol{s}$ are the occupations of the different Slater determinants. Here $|\Psi_{c}^{t}\rangle$ is a singly excited Slater determinant, where the occupied spin-orbital $c$ is excited to the $t$-th unoccupied orbital. The usual orthogonality conditions between the excited Slater-determinants hold $\langle\Psi_{c}^{t}|\Psi_{d}^{u}\rangle=\delta_{c,d}\delta_{t,u}$ and like-wise with respect to $|\Psi_{0}\rangle$50 To simplify our discussion further, we restrict our considerations to closed-shell cHF in the following, i.e., the spin orbital $\phi_{i}(\boldsymbol{\tau})$ representation of the electron integrals turn into a spatial orbital representation $\chi_{i}(\boldsymbol{r})$. In particular, every orbital is assumed to be doubly occupied by electrons. To simplify our notation, the spatial DSE and Coulomb two-electron integrals are given as follows:
$(ud|ct)_{\rm DSE}$ $=$ $\langle\chi_{u}|\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}}|\chi_{d}\rangle\langle\chi_{c}|\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}}|\chi_{t}\rangle,$ (20)
[…]
with $(ud|ct)=(ud|ct)_{\rm C}+(ud|ct)_{\rm DSE}$. We note that the standard Brillouin theorem $\langle\Psi_{0}|\hat{H}^{e}|\Psi_{c}^{t}\rangle=0$, as well as the Slater-Condon rules remain valid for the DSE two-electron integrals. Consequently, the transition matrix element between singlet symmetry-adapted configurations from real orbitals reduce to the following compact form 50
$\langle\Psi_{d}^{u}|\hat{H}^{e}|\Psi_{c}^{t}\rangle=(E_{0}+\epsilon_{t}-\epsilon_{c})\delta_{d,c}\delta_{u,t}+2(ud|ct)-(ut|cd).$ (22)
The resulting CI-singles energies are
$\mathscr{E}_{\boldsymbol{s}}$ $=\langle\Phi_{\boldsymbol{s}}|\hat{H}^{e}|\Phi_{\boldsymbol{s}}\rangle$ (23)
$=\frac{1}{\sum_{ev}(s_{e}^{v})^{2}}\sum_{c}^{N}\sum_{d}^{N}\sum_{t}^{\infty}\sum_{u}^{\infty}\bigg{(}s_{c}^{t}s_{d}^{u}\big{[}(E_{0}+\epsilon_{t}-\epsilon_{c})\delta_{d,c}\delta_{u,t}+2(ud|ct)-(ut|cd)\big{]}\bigg{)}.$
By ordering all the CI-singles energies we can then explore the lowest lying excitations and also potential degeneracies.
To again make things a little simpler and focus on the essentials, i.e., the long-range interactions induced by the cavity, we make a quasi-dilute gas assumption. That is, we will partition our orbitals $c,d,t,u$ into a set $S_{\rm inter}$, for which we have
$2(ud|ct)_{\rm C}-(ut|cd)_{\rm C}\rightarrow 0\quad{\rm yet}\quad 2(ud|ct)_{\rm DSE}-(ut|cd)_{\rm DSE}\neq 0$ (24)
At the same time we assume that the rest of the orbitals are in $S_{\rm intra}$, for which
$2(ud|ct)_{\rm C}-(ut|cd)_{\rm C}\neq 0\quad{\rm yet}\quad 2(ud|ct)_{\rm DSE}-(ut|cd)_{\rm DSE}\rightarrow 0$ (25)
In other words, we assume a partitioning of our set of orbitals into intramolecular orbitals in $S_{\rm intra}$, which are the focus of standard (single-molecule) quantum chemistry, and into intermolecular orbitals in $S_{\rm inter}$, which are de-localized orbitals. For the latter, the long-range properties of the transverse DSE interaction start to dominate over the short-ranged (longitudinal) Coulomb interaction. For localized molecular orbitals, which determine the intramolecular structure, the Coulomb two-electron integrals will be decisive and thus cannot be discarded. Thus Eq. (23) can be separated as
$\mathscr{E}_{\boldsymbol{s}}$ $=$ $E_{0}+E_{\boldsymbol{s}}^{\rm C}+E^{\rm DSE}_{\boldsymbol{s}}$ (26)
$\approx$ $E_{0}+E^{\rm C,intra}_{\boldsymbol{s}}+E^{\rm C,inter}_{\boldsymbol{s}}+E^{\rm DSE,inter}_{\boldsymbol{s}}$
$\overset{\rm quasi\ dilute}{\approx}$ $E_{0}+E^{\rm C,intra}_{\boldsymbol{s}}+E^{\rm DSE,inter}_{\boldsymbol{s}},$
where the first approximation is a consequence of the typically small contributions of the DSE interaction for localized orbitals, whereas the second approximation is only reasonable, if the quasi-dilute gas approximation applies for all molecules within our ensemble of interest.
Using the quasi-dilute gas picture in Eq. (26), we can consider the DSE correlation energy $E^{\rm DSE,inter}_{\boldsymbol{s}}=\big{(}\sum_{c,d,t,u\ \in S_{\rm inter}}s_{c}^{t}s_{d}^{u}\big{[}(\epsilon_{t}-\epsilon_{c})\delta_{d,c}\delta_{u,t}+2(ud|ct)-(ut|cd)\big{]}\big{)}/\big{(}\sum_{ev}(s_{e}^{v})^{2}\big{)}$ independent from the local intramolecular correlation energy described by $E^{\rm C,intra}_{\boldsymbol{s}}=\big{(}\sum_{c,d,t,u\ \in S_{\rm intra}}s_{c}^{t}s_{d}^{u}\big{[}(\epsilon_{t}-\epsilon_{c})\delta_{d,c}\delta_{u,t}+2(ud|ct)-(ut|cd)\big{]}\big{)}/\big{(}\sum_{ev}(s_{e}^{v})^{2}\big{)}$. That is, the intra- and the intermolecular energies decouple and we can exclusively focus on the intramolecular energy contributions. In a next step we note that
$\sum_{c}^{N}\sum_{d}^{N}\sum_{t}^{\infty}\sum_{u}^{\infty}s_{c}^{t}s_{d}^{u}(ud|ct)_{\rm DSE}=\bigg{(}\sum_{c,t}s_{c}^{t}\langle\chi_{c}|\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}}|\chi_{t}\rangle\bigg{)}^{2}\geq 0.$ (27)
That is, the correlations have a similar bi-partite Hartree term as the cHF equation for $|\Psi_{0}\rangle$. This allows to restructure the DSE correlation energy as
$E^{\rm DSE,intra}_{\boldsymbol{s}}=$ $\frac{1}{\sum_{ev}(s_{e}^{v})^{2}}\bigg{[}\sum_{c,t}(s_{c}^{t})^{2}\underset{=-\mathscr{K}_{cc}^{tt}}{\underbrace{\big{(}\epsilon_{t}-\epsilon_{c}-(tt|cc)_{\rm DSE}\big{)}}}-\sum_{c,d}\sum_{t,u}s_{c}^{t}s_{d}^{u}\underset{\mathscr{K}_{cd}^{tu}}{\underbrace{(ut|cd)_{\rm DSE}(1-\delta_{cd}\delta_{tu})}}$
$+\bigg{(}\sum_{c,t}s_{c}^{t}\langle\chi_{c}|\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}}|\chi_{t}\rangle\bigg{)}^{2}$ (28)
$=-\sum_{c,d}\sum_{t,u}s_{d}^{u}s_{c}^{t}\frac{\mathscr{K}_{cd}^{tu}}{\sum_{ev}(s_{e}^{v})^{2}}+\frac{2}{\sum_{ev}(s_{e}^{v})^{2}}\bigg{(}\sum_{c,t}s_{c}^{t}\langle\chi_{c}|\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}}|\chi_{t}\rangle\bigg{)}^{2}$ (29)
Since the second term of Eq. (29) is positive and thus its minimal value is zero, when searching for the lowest correlation energies, this term can be incorporated as a side condition that fixes one occupation number $s_{d}^{u}$. That is, we can perform a minimzation by enforcing $\sum_{c,t}s_{c}^{t}\langle\chi_{c}|\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}}|\chi_{t}\rangle=0$. We will thus discard this energy contribution and only take it into account by restricting the space of allowed states to vary over. In a next step we re-label the orbital occupations by
$s_{i}=\frac{s_{c}^{t}}{\sqrt{\sum_{e,v}(s_{e}^{v})^{2}}},$ (30)
and the overlaps by
$\mathscr{J}_{ij}=\frac{\mathscr{K}_{cd}^{tu}}{\sum_{ev}(s_{e}^{v})^{2}},$ (31)
such that we end up with
$E^{\rm DSE,intra}_{\boldsymbol{s}}=-\sum_{i,j}s_{i}s_{j}\mathscr{J}_{ij}.$ (32)
For sufficiently large molecular ensembles within the quasi-dilute gas regime, the resulting $\mathscr{J}_{ij}$ can be regarded as a random variable. Its distribution depends on the respective orbital excitation energy $\epsilon_{t}-\epsilon_{c}\mapsto\Delta\epsilon_{i}$, as well as on the orientation (polarization) of the the excitation with respect to the polarization of the relevant cavity modes, (diagonal and off-diagonal contributions from $(ut|cd)_{\rm DSE}$-terms, see Eq. (28)). As a consequence, Eq. (32) is similar to a spin glass Hamiltonian. 108, 109 The probability distribution of $\mathscr{J}_{ij}$ will strongly depend on the molecular properties of the collectively coupled ensemble. In more detail, to expect non-negligible DSE correlation effects (i.e., $\exists\;i\neq j,\ {\rm with}\ |\mathscr{J}_{ij}|>0$), very de-localized electronic orbitals (excitations) are required. However, most of this de-localized excitations will also possess a high orbital excitation energy $\Delta\epsilon_{i}$ and can thus be safely discarded from the considerations in Eq. (32) when exploring the lowest electronic energy landscape. Consequently, to expect significant intermolecular DSE correlation effects, we additionally require a highly (almost) degenerate HF ground state (i.e. $\exists i,\ {\rm with}\ \langle\mathscr{J}_{ii}\rangle\approx 0$). For sufficiently strong collective coupling, we have already seen that such a highly degenerate electronic ground state (polarization glass) exists in the dilute gas limit (see Sec. 2.3.1). Increasing slightly the intermolecular densities, i.e., going to the quasi-dilute limit, we expect that at least for specific chemical setups we find a non-vanishing intermolecular DSE correlation energy of the following form
$E_{\rm corr}^{\rm DSE}=-\sum_{i<j}^{N_{J}}s_{i}s_{j}J_{ij},\ \sum_{i}s_{i}^{2}=1,\ \langle J_{ii}\rangle=0,$ (33)
where $N_{J}$ is the number of the relevant (almost zero energy when compared to the relevant scale) orbital excitations in the ensemble. We expect the distribution of the random variables $J_{ij}$ to be a heavily-tailed distribution (e.g. Cauchy-like), since most excitations will contribute only a little, but there might be a few very de-localized excitations, which contribute significantly. Understanding the spin glass properties of Eq. (33) for such distributions will require non-trivial and computationally expensive simulation. Finally, we note that a cavity-mediated spin glass phase is not an entirely new concept. However, previous theoretical concepts as well as experimental realizations rely on the complex restructuring of the photon modes to reach the desired random interactions.110, 111, 112, 113, 114, 115, 116 Instead, the here proposed concept utilizes the complex electronic-structure of molecules to create a cavity-mediated spin glass. In more detail, the origin of randomness emerges from the interplay between breaking the isotropy of space and orienting the molecules randomly with respect to the distinguished polarization axis.
4 The physics of a spin glass
In the previous sections we have established a formal connection between the physics of spin glasses and the cavity-induced inter-molecular electron correlations under VSC. Thanks to the theoretical similarities between both problems, we are confident that the polaritonic-chemistry community can learn from established knowledge of the spin glass community. In the following, we briefly introduce some key-concepts of spin glasses using the Sherrington-Kirkpatrick (SK) model. It provides a paradigmatic model of a spin glass with long-range (all-to-all) interaction, for which exact results have been determined after decades of intensive research that ultimately were awarded with the Nobel prize in physics (2021). 117 We note that many models of spin glasses are based on (random) short-ranged interactions between spins on a crystal lattice, e.g., the Edwards-Anderson model 88. Furthermore, the classical spin variable are usually discrete. However, in a cavity we have a continuous long-range (all-to-all) electron correlation interaction. The spin glass properties of the SK model are considered generic, since it serves as a limiting case 118 for the Bethe lattice model, 119 the long-range Edwards-Anderson model and the infinite dimensional Edwards-Anderson model. 88. Those key concepts, for example, involve spin glass phase transition, frustration, replica symmetry breaking, (off)-equilibrium fluctuations and aging effects. Eventually, after having introduced above concepts for the SK model, we will have a closer look at the spherical 2-spin glass model,109 that provides (to our knowledge) the closest known spin glass model to describe the cavity-mediated electron correlations derived in Eq. (33).
Figure 4: Pictorial representation of a cavity-mediated molecular polarization/spin glass under by vibrational strong coupling. White arrows indicate the locally modified (slighly polarized) electronic-structure. Figure taken from Ref.20.
4.1 The Sherrington-Kirkpatrick (SK) model
@@
$H_{SK}(\boldsymbol{\sigma})=-\sum_{j<i}^{N_{J}}J_{ij}\sigma_{i}\sigma_{j},\ ,\ J\sim\mathscr{N}(J_{0}/N_{J},\tilde{J}^{2}/N_{J})$ (34)
and the couplings $J_{ij}$ between the spins are independent random variables that are normally distributed. The spins take discrete values $\sigma_{i}\in\{\pm 1\}$. The presence of an additional external magnetization field $h$, acting on the spins as $\sum_{i}^{N_{J}}h\sigma_{i}$ to generate a finite magnetization $m$, can be recast into a finite mean-value $J_{0}=h/m$ of the random interactions $J_{ij}$120 To understand the basic physical properties of the SK model at finite temperature $T$, we follow Refs. 117, 118 and continue with some definitions. The local (”single molecule”) magnetization (in the case of polaritonics this would be polarization) at temperature $T$ is given by,
@@
where $\alpha$ denotes a possible quasi-thermal equilibrium state of the SK model, i.e., a local minimum in the phase space for a given choice of $J$, at temperature $T$. In more detail, the magnetizations $m(i)_{\alpha}$ correspond to the $\alpha$-th solution of the mean-field equation, 121, 118
@@
which becomes an exact description for the SK model in the thermodynamic limit. 89, 122, 118 The determination of the exponentially large number of solutions is a computationally very demanding task. Eventually, the thermal equilibrium (”single molecular”) magnetization, averaged over all $\alpha$ and all possible choices of J, is defined as
@@
We note that averaging over all $\alpha$ and $J$ makes the result site-independent, i.e., the left-hand side of Eq. (37) is independent of $i$. The numerous fundamental physical properties of the SK model and its implications the understanding of spin glasses in general will be discussed in the following with its implications on polaritonic chemistry (see Sec. 5).
4.1.1 Spin glass phase
To determine the phase diagram of the SK model, we introduce the Edwards-Anderson order parameter or self-overlap as defined in Ref. 88
which can be shown to neither depend on the state $\alpha$ nor the specific realization of J.118, 117 Based on the magnetization $m$ and the magnetic order parameter $q_{EA}$, Sherrington and Kirkpatrick determined an analytical phase diagram in the thermodynamic limit $N\rightarrow\infty$. Their computations suggested the emergence of three different phases: a ferromagnetic ($q_{EA}\neq 0$, $m\neq 0$), a spin glass ($q_{EA}\neq 0$, $m=0$) and a paramagnetic one ($q_{EA}=0$, $m=0$). However, Sherrington and Kirkpatrick already noticed that their solution seemed questionable at low temperatures, since it gave rise to a negative entropy, which is nonphysical.89 Indeed, Almeida and Thouless showed that the original solution of the SK model becomes unstable in the thermodynamic limit at sufficiently low temperature, which leads to the corrected phase diagram of the SK model displayed in Fig. 5120 Almeida and Thouless could determine an explicit stability criterion at low temperature for the SK model 120
$k_{B}T>\frac{4}{3\sqrt{2\pi}}Je^{-\frac{J_{0}^{2}}{2J^{2}}}.$ (39)
The striking feature of this result is that in the ground state ($T\rightarrow 0$), the solution of the SK model becomes unstable, and thus enters the spin glass phase, even if $\infty>J_{0}\gg J$, i.e., even if the system is exposed to very strong external magnetization fields.
For a polaritonic setting this could imply that even a (very) small intermolecular DSE correlation interaction can potentially introduce a spin glass-type phase transition that fundamentally alters microscopic and macroscopic properties of a polaritonic ensemble at low-enough temperatures. The SK analogy also highlights that the temperature for the electronic subsystem might become important when treating the dressed electronic-structure under VSC. Even though we assumed bare molecules, for which non-adiabatic coupling effects to the intramolecular excited electronic-structures are irrelevant in free space.
The origin of the Almeida and Thouless instability can be attributed to spontaneous replica symmetry breaking that we briefly discuss next.117 Notice further, that a spin glass is conceptually distinct from the Anderson localization mechanism, which has been discussed in the context of polaritonic transport properties. 123, 124
Figure 5: Phase diagram of the SK model, reproduced from Ref. 120. The doted line corresponds to the original (erroneous) spin glass phase proposed by Sherrington and Kirkpatrick, which did not consider that spontaneous replica symmetry breaking can occur. Replica symmetry breaking extends the (unstable) spin glass regime at low temperature to much larger $J_{0}/J\gg 0$ values (dark red). Notably, in the ground state ($T\rightarrow 0$), the instability occurs for arbitrarily large (but finite) $J_{0}$-values, provided that $J>0$.
4.1.2 Free energy and spontaneous replica symmetry breaking
In order to better understand physical consequences of the SK model defined in Eq. (34), we have a closer look at the average free energy $F$, which can be defined as follows in the thermodynamic limit $N\rightarrow\infty$117
$F(T)$ $=$ $-\lim_{N\rightarrow\infty}\frac{\overline{k_{B}T\log(Z_{J}(T,N))}}{N},$ (40)
$Z_{J}(T,N)$ $=$ $2^{-N}\sum_{\{\boldsymbol{\sigma}\}}e^{-\frac{H_{J}(\boldsymbol{\sigma})}{k_{B}T}}.$ (41)
The overline indicates the averaging of the partition function $Z_{J}$ with respect to randomly drawn $J_{ij}$ realizations. To simplify the $J$-averaging, the so-called replica trick was proposed, where instead of one system with $N$-spins, an extended system consisting of $n$-times the same system made of $N$-spins is considered. It simplifies taking the logarithm as follows 117
$F_{n}(T)$ $=$ $-\lim_{N\rightarrow\infty}\frac{\overline{k_{B}T(Z_{J}(T,N))^{n}}}{nN},$ (42)
$F(T)$ $=$ $\lim_{n\rightarrow 0}F_{n}(T).$ (43)
The replica trick works if $F_{n}$ is analytic in $n$ and has no singularities. 117 In particular, one would expect replica symmetry to hold as a natural assumption, since it implies the re-shuffling of the $n$ identical replicas will not change the result. From this symmetry aspect, one can deduce that the free energy depends on a single order parameter $q$, which then can be minimized to determine the alleged solution of the SK model. 89, 117 However, it turns out that $F_{n}$ is indeed not analytic for $n<n_{c}<1$ in the SK model. This indicates that the replica symmetry is spontaneously broken and thus things become much more complex. After a long endeavour the exact solution of the replica Ansatz was discovered by Parisi yielding the following free energy 122, 125
$F$ $=$ $\max_{\mathscr{q}(\mathscr{x})}F[\mathscr{q}(\mathscr{x})],$ (44)